OpenStudy (anonymous):

In a certain county, the number of Charter Schools is six less than twice the number of Alternative Schools. We know that there are 52 Charter Schools in the county. How many Alternative Schools are in the county? a. 58 b. 44 c. 29 d. 23

OpenStudy (anonymous):

I believe the answer is C, 29

OpenStudy (anonymous):

Alternative schools = x \[2x-6=52\] This means 2 * Alternative schools - 6 = Charter schools \[2x-6+6=52+6\] -6+6 cancel out \[2x=52+6\] Add up \[2x=58\] Divide by 2 \[ \frac{ 2x }{ 2 }=\frac{ 58 }{ 2 }\] \[x=29\] There are 29 Alternative schools.

OpenStudy (anonymous):

Alternative schools =x and the number of charter schools = 2x-6. Since there are are 52 charter schools we can say that 52=2x-6, now solve for x. u plug in 29 and it is 29

OpenStudy (anonymous):

you guys are awsome :)